### v1

@@ Abstract
Context. The limited number of high-resolution spectra of hot stars is inadequate for statistical studies of DIBs. In contrast, the vast quantity of low-resolution spectroscopic survey on cool stars holds great potential for investigating the relationship between DIBs and known interstellar medium (ISM), as well as the spatial distribution of their unidentified carriers.
Aims. We attempt to measure the DIBs $\lambda 5780$, $\lambda 5797$ and $\lambda 6614$ in over two million low-resolution spectra of cool stars from LAMOST. Based on the DIB measurements, the correlation between DIBs and extinction, the kinematics of DIBs, and the Galactic distribution of DIBs are reviewed and investigated from the perspective of statistics.
Methods. A pipeline is developed to measure the DIBs $\lambda 5780$, $\lambda 5797$ and $\lambda 6614$ in the LAMOST low-resolution spectra. Four modules in the pipeline consist of building the target and reference dataset, extracting the ISM residual spectra from the target spectra, measuring the DIBs in the residual spectra, and quality control of the measurements.
Results. We obtain the DIB measurements of spectra of late-type stars from LAMOST, and screen out 176,831, 13,473 and 110,152 high-quality (HQ) measurements of the DIBs $\lambda 5780$, $\lambda 5797$ and $\lambda 6614$, respectively, corresponding to 142,074, 11,480 and 85,301 unique sources. Utilizing these HQ measurements, we present the Galactic maps of the DIBs $\lambda 5780$ and $\lambda 6614$ in the northern sky for the first time. The central wavelengths of the DIBs $\lambda 5780$, $\lambda 5797$ and $\lambda 6614$ in air are determined to be $5780.48\pm 0.01$, $5796.94\pm 0.02$ and $6613.64\pm 0.01$, respectively, based on their kinematics. The equivalent widths of these three DIBs per unit extinction are statistically fitted to be 0.565, 0.176 and 0.256 Å$\rm\AA\,mag^{-1}$. As a part of our work, three catalogs of the HQ measurements for the DIBs $\lambda 5780$, $\lambda 5797$ and $\lambda 6614$ are provided via https://nadc.china-vo.org/res/r101404/.
Conclusions. To the best of our knowledge, this is the largest number of measurements of these three DIBs to date. It is also the first time that the Galactic maps of the DIBs $\lambda 5780$ and $\lambda 6614$ in the northern hemisphere are presented, and the central wavelengths of the DIBs $\lambda 5780$, $\lambda 5797$ and $\lambda 6614$ are estimated from the kinematics.
@@ 1 Introduction
Diffuse interstellar bands (DIBs) are a set of broad interstellar absorption features that are ubiquitously observed from the optical to the near-infrared wavelength range (Fan et al. 2019; Ebenbichler et al. 2022; Hamano et al. 2022; Vogrinčič et al. 2023; Castellanos et al. 2024). Since the first discovery of the DIBs $\lambda 5780$ and $\lambda 5797$  [^1] We follow a naming convention based on the approximate central wavelength (Å) of DIBs in air. by Heger (1922), plentiful efforts have been dedicated to identifying the carriers of DIBs (e.g., Iglesias-Groth 2007; Krełowski et al. 2011; Zanolli et al. 2023), but the question remains open, except $\rm C_{60}^{+}$ (Campbell et al. 2015) that can be responsible for five DIBs in the near-infrared wavelength range (see a review by Linnartz et al. 2020).
@@
To bind the DIB carriers and guide the laboratory search for them, attention has been turned to the astronomical observational facts, such as their own profiles and their correlations with the other tracers of interstellar medium (ISM). For example, MacIsaac et al. (2022) constrained the sizes of possible carrier molecules by comparing the peak-to-peak separation in the DIBs $\lambda 5797$, $\lambda 6379$ and $\lambda 6614$. To better understand the behavior of DIBs in different interstellar environments, Fan et al. (2017) analysed how the pattern of the equivalent widths (EWs) of DIB normalized by the corresponding extinction changes with the molecular fraction $f_{\rm H2}$.
@@
As the strongest and earliest observed DIBs (Merrill 1930), the DIBs $\lambda 5780$, $\lambda 5797$ and $\lambda 6614$, as well as their interrelations have been extensively studied and discussed in the literature. Owing to the close proximity of the DIBs $\lambda 5780$ and $\lambda 5797$, they are usually treated as a pair to investigate. The well-known “$\sigma-\zeta$” effect was found from the varying strength ratio of these two prominent DIBs (e.g., Krelowski & Westerlund 1988; Vos et al. 2011; Kos & Zwitter 2013). These works indicated that the EWs of the DIB $\lambda 5780$ in the “$\zeta-$type” sight-lines tend to be weaker than those in the “$\sigma-$type” sight-lines while the corresponding EWs of the DIB $\lambda 5797$ stay unchanged. Upon discovering that the relationship between DIB and reddening is linear but not excellent (including the DIBs $\lambda 5780$, $\lambda 5797$ and $\lambda 6614$; Raimond et al. 2012; Puspitarini et al. 2013), and there exist the “skin effect” (Snow & Cohen 1974; Herbig 1995), the correlations with the other ISM tracers, e.g., OH, CH, CO, H  I and $\rm H_{2}$, have been explored (Weselak et al. 2004; Friedman et al. 2011; Lan et al. 2015; Fan et al. 2017; Weselak 2019). It is now well established from those studies that the DIB $\lambda 5797$ is more closely related to column of density of $\rm H_{2}$ while the DIB $\lambda 5780$ to that of H  I . In addition, through the studies on the correlation between DIBs, a nearly perfectly correlated DIB pair, i.e., DIB $\lambda 6196$-$\lambda 6614$ (Cami et al. 1997; Moutou et al. 1999; Galazutdinov et al. 2002; McCall et al. 2010; Bailey et al. 2016), some decently correlated DIB pairs, e.g., DIB $\lambda 5780$-$\lambda 6614$ and DIB $\lambda 5797$-$\lambda 5850$ (Friedman et al. 2011; Smith et al. 2021; Fan et al. 2022), and even some anti-correlated DIB pairs, e.g., DIB $\lambda 4984$-$\lambda 7559$ and $\lambda 5418$-$\lambda 7562$ (Fan et al. 2022), have been identified, which can help to group the DIBs into specific DIB families. Furthermore, in high-resolution spectra, the longward blue wing in the DIB $\lambda 5780$ profile (Snell & vanden Bout 1981; Danks & Lambert 1976), the asymmetry profile in the DIB $\lambda 5797$ (MacIsaac et al. 2022), and the triple-peak fine structure in the DIB $\lambda 6614$ profile (Herbig & Soderblom 1982; Galazutdinov et al. 2008), all of which point their carriers to the complex macromolecules.
@@
However, the majority of the aforementioned studies are mainly based on the high-resolution spectra of early-type stars without contamination of the stellar lines. It is challenging to carry out larger-scale and more statistically significant research as a result of the limited sight-lines towards the early-type stars. Kos et al. (2013) introduced a best neighbor matching method (BNM) to extract the DIB $\lambda 8621$ from $\sim$ 500,000 spectra of late-type stars in RAVE (Steinmetz et al. 2006) and provided a statistical ratio of the DIB $\lambda 8621$ to the extinction. Although the objects observed in SDSS (Yanny et al. 2009) do not concentrate on the Galactic plane where the reddening is high, Lan et al. (2015) still managed to measure over 20 DIBs from the spectra of cool stars by stacking the spectra with similar extinction, and map out the distribution of DIB at high latitudes in the Milky Way. Later on, using the cool stars from APOGEE (Majewski et al. 2017), Zasowski et al. (2015) presented the projected Galactic map of the single DIB at 1.5273 $\mu$m and firstly estimated its precise central wavelength by the kinematics information. Following the same idea, Gaia Collaboration et al. (2023b) and Gaia Collaboration et al. (2023a) elaborated the properties of the DIBs $\lambda 8621$ and $\lambda 8648$ in the $Gaia$ RVS spectra (Cropper et al. 2018; Seabroke et al. 2021) of both cool and hot stars. Furthermore, Cox et al. (2024) constructed the 3D Galactic map traced by the DIB $\lambda 8621$ based on the measurements by Gaia Collaboration et al. (2023b). Benefiting from 872,000 cool stars in GALAH (Buder et al. 2021), brand new DIBs were found in residuals of these stellar spectra (Vogrinčič et al. 2023). Moreover, Saydjari et al. (2023) and Zhao et al. (2024) have recently employed the method based on machine learning to improve the measurement of DIB from the spectra of cool stars. These studies on DIB measurement from cool stellar spectra all involve removing the stellar lines from spectra to extract the ISM composition. The critical distinction between these methods lies in how the stellar lines are obtained. On the one hand, studies such as Chen et al. (2013), Puspitarini et al. (2015), Zhao et al. (2021a) and Gaia Collaboration et al. (2023b) use theoretical methods to synthesize the stellar lines. On the other hand, the observed spectra with low or no extinction are utilized to pairwise match the stellar components (e.g., Cordiner et al. 2008a, b; Kos & Zwitter 2013; Kos et al. 2014; Zasowski et al. 2015 and Gaia Collaboration et al. 2023a), which is also the approach adopted in this study.
@@
Besides the works of Yuan & Liu (2012) and Lan et al. (2015) which are based on the SDSS low-resolution spectra (R $\approx$ 2000), certain teams have also made efforts to dig out treasure from low-resolution spectra. Using the spectra with a resolving power R = 3300, Cordiner et al. (2008a) investigated the DIBs $\lambda 5780$, $\lambda 5797$, $\lambda 6203$, $\lambda 6283$ and $\lambda 6614$ in the ISM of M31. Then, they still used the spectra with similar resolution to extend their work to the DIBs in the ISM of M33 (Cordiner et al. 2008b), and undertook the first survey of DIBs observed in the spectra of B-type supergiants in M31 (Cordiner et al. 2011). By means of the DIBs $\lambda 5780$ and $\lambda 5797$ measured from the spectra (Farhang et al. 2015a) with R = $\sim$ 2000 in the northern hemisphere, Farhang et al. (2015b) probed the Local Bubble and its surroundings. Later, combining the DIB measurements derived from the spectra with a resolution of 5500 in the southern observations, they presented a 3D map of the hot Local Bubble (Farhang et al. 2019). Although the quantity of low-resolution spectra used in these studies has seen a tremendous leap compared to high-resolution spectra, it is still insufficient to support a broader perspective for investigating the properties of DIBs, e.g., their spatial distribution and kinematics in the Milky Way.
@@
In this work, we take full advantages of the LAMOST low-resolution survey (LAMOST LRS; Zhao et al. 2012; Deng et al. 2012; Luo et al. 2015). The survey almost covers the entire northern sky from the declination of -10 to nearly 90, especially the Galactic plane where ISM is rich. So far, there have been a few millions of spectra of cool stars across the optical wavelength range, which allow us to detect several DIBs simultaneously from the perspective of big data mining. The DIBs $\lambda 5780$, $\lambda 5797$ and $\lambda 6614$ are selected as our subjects due to their higher intensities compared to the other DIBs (Fan et al. 2019; Vogrinčič et al. 2023), which are more likely to be straightforward measured in the individual spectrum of cool star. Although the DIBs $\lambda 4430$ and $\lambda 6283$ are also strong, they are not included in this work because of the complexity of the spectral region nearby the DIBs $\lambda 4430$ and $\lambda 6283$. For the DIB $\lambda 4430$, it is hard to determine the continuum level due to the adjacent strong hydrogen Balmer lines and its own broad profile. For the DIB $\lambda 6283$, the O2 telluric band wraps around it (Herbig 1975; Jenniskens & Desert 1994) and is not cleanly removed in the LAMOST spectra. As a sequence, we focus on the DIBs $\lambda 5780$, $\lambda 5797$ and $\lambda 6614$, build the largest sample of measurements of these three DIBs to date, and conduct a comprehensive analysis of their properties in the context of big data.
@@
The remainder of this paper is organized as follows. In Section 2, we describe the data used in this work. The pipeline for measuring the DIBs $\lambda 5780$, $\lambda 5797$ and $\lambda 6614$ in spectra of cool stars is detailed in Section 3. In Section 4, we perform quality control and a series of validation tests to ensure the reliability of the measurements. The results and discussions of the measurements are presented in Section 5. Section 6 introduces the code and data availability. Finally, we summarize our findings in Section 7.
@@ 2 Data
LAMOST (the Large Sky Area Multi-Object-fiber Spectroscopic Telescope; Zhao et al. 2012; Deng et al. 2012; Luo et al. 2015) is a reflecting Schmidt telescope with a 5° field-of-view and an effective aperture of 4 meters. Thanks to 4000 fibers averagely distributed among 16 spectrometers (i.e. each one accepts 250 fibers), LAMOST is able to simultaneously observe 4000 objects in theory. There are two resolving modes for LAMOST: one is the low resolution (R = 1800 at 5500 Å) and the other is the medium resolution (R = 7500 at 5163 Å and 6593 Å). When switching to the low-resolution mode employed here, the designed optical band covers the range from 3700 Å to 9000 Å.
The data used in this work are from the data release 10 of LAMOST low-resolution spectroscopic survey (LAMOST LRS DR10)  [^2] https://www.lamost.org/dr10.. A total of 11,817,430 spectra are published in this release, which contains 11,473,644 stellar spectra, 263,444 galaxy spectra, and 80,342 quasar spectra. All the published spectra are reduced by wavelength calibration, sky background subtraction, telluric correction and relative flux calibration (Zhao et al. 2012; Luo et al. 2015). Note that the telluric absorption is negligible in the regions of the DIBs $\lambda 5780$, $\lambda 5797$ and $\lambda 6614$ investigated in this work (see Figure 16 in Matheson et al. 2000 and Figure 2 in Kimeswenger et al. 2015 for more details). All the LAMOST targets are cross-matched with $Gaia$ DR3 sources (Gaia Collaboration et al. 2023c) within a radius of 3 arcsec, and the $Gaia$ IDs, i.e., gaia_source_id, are included in the official LAMOST DR10 catalog  [^3] https://www.lamost.org/dr10/v1.0/catalogue. The stellar spectra we focus on are corresponding to $\sim$2,000,000 unique stars and the majority of them concentrate on the Galactic plane. For the cool stars, i.e., the F, G, K and late A-type stars, the LAMOST stellar parameter pipeline (LASP) provides the effective temperature ($T_{\rm eff}$), surface gravity ($\log g$), metallicity abundance ([Fe/H]), and radial velocity (RV) with the precision of 100 K, 0.19 dex, 0.14 dex, and 6 $\rm km\ s^{-1}$, respectively  [^4] Here we refer reader to the note of LAMOST DR10 via https://www.lamost.org/dr10/v1.0/doc/release-note.. In addition, the continuum-normalized flux are calculated for these spectra with stellar parameters (Song et al. 2012).
Benefiting from the estimations of geometric distances from parallaxes in $Gaia$ DR3 (Bailer-Jones et al. 2021), the relatively reliable distance of an object in LAMOST can be simply derived. Combined with the three-dimensional dust reddening map (version name “bayestar2019”, Green et al. 2019), once the celestial coordinate and astrometric distance of an object is given, the fine-grained extinction $\rm E(B-V)$  [^5] A recalibration factor of 0.884 (Schlafly & Finkbeiner 2011) is applied for $\rm E(B-V)$. can be calculated at the sight-line of each object.

### v2

@@ Abstract
Context. The limited number of high-resolution spectra of hot stars is inadequate for statistical studies of diffuse interstellar bands (DIBs). In contrast, the vast quantity of low-resolution spectroscopic surveys on cool stars holds great potential for investigating the relationship between DIBs and the known interstellar medium (ISM), as well as the spatial distribution of their unidentified carriers.
Aims. We attempt to measure the DIBs $\lambda 5780$, $\lambda 5797$, and $\lambda 6614$ in over two million low-resolution spectra of cool stars from the Large Sky Area Multi-Object Fiber Spectroscopic Telescope (LAMOST). Based on these DIB measurements, we reviewed and investigated the correlation between DIBs and extinction; the kinematics of DIBs; and the Galactic distribution of DIBs from a statistical perspective.
Methods. We developed a pipeline to measure the DIBs $\lambda 5780$, $\lambda 5797$, and $\lambda 6614$ in the LAMOST low-resolution spectra. Four modules in the pipeline consist of building the target and reference dataset; extracting the ISM residual spectra from the target spectra; measuring the DIBs in the residual spectra; and quality control of the measurements.
Results. We obtained DIB measurements of spectra of late-type stars from LAMOST, and selected 176,831, 13,473, and 110,152 high-quality (HQ) measurements of the DIBs $\lambda 5780$, $\lambda 5797$, and $\lambda 6614$, respectively, corresponding to 142,074, 11,480, and 85,301 unique sources. Using these HQ measurements, we present Galactic maps of the DIBs $\lambda 5780$ and $\lambda 6614$ in the northern sky for the first time. The central wavelengths of the DIBs $\lambda 5780$, $\lambda 5797$, and $\lambda 6614$ in air are determined to be $5780.48\pm 0.01$, $5796.94\pm 0.02$, and $6613.64\pm 0.01$, respectively, based on their kinematics. A statistical fit of the equivalent widths of these three DIBs per unit extinction provides values of 0.565, 0.176, and 0.256 Å$\rm\AA\,mag^{-1}$. As a result of this work, three catalogs of the HQ measurements for the DIBs $\lambda 5780$, $\lambda 5797$, and $\lambda 6614$ are provided via https://nadc.china-vo.org/res/r101404/.
Conclusions. To the best of our knowledge, this is the largest number of measurements of these three DIBs to date. It is also the first time that Galactic maps of the DIBs $\lambda 5780$ and $\lambda 6614$ in the northern hemisphere are presented, and that the central wavelengths of the DIBs $\lambda 5780$, $\lambda 5797$, and $\lambda 6614$ are estimated from kinematics.
@@ 1 Introduction
Diffuse interstellar bands (DIBs) are a set of broad interstellar absorption features that are ubiquitously observed from the optical to the near-infrared (NIR) wavelength range (Fan et al. 2019; Ebenbichler et al. 2022; Hamano et al. 2022; Vogrinčič et al. 2023; Castellanos et al. 2024). Since the first discovery of the DIBs $\lambda 5780$ and $\lambda 5797$  [^1] We follow a naming convention based on the approximate central wavelength (Å) of DIBs in air. by Heger (1922), many efforts have been dedicated to identifying the carriers of DIBs (e.g., Iglesias-Groth 2007; Krełowski et al. 2011; Zanolli et al. 2023), but with the exception of $\rm C_{60}^{+}$ (Campbell et al. 2015), which has been found to be responsible for five DIBs in the NIR wavelength range (see a review by Linnartz et al. 2020), the carriers of DIBs remain unknown.
@@
In attempts to identify the carriers of DIBs and guide laboratory searches for them, attention has been turned to the characterization of DIBs based on astronomical observations, such as their intrinsic profiles in high-resolution spectra and their correlations with other tracers of the interstellar medium (ISM). For example, MacIsaac et al. (2022) constrained the sizes of possible carrier molecules by comparing the peak-to-peak separation in the DIBs $\lambda 5797$, $\lambda 6379$, and $\lambda 6614$ using the extensive spectra from the ESO diffuse interstellar bands large exploration survey (EDIBLES). To better understand the behavior of DIBs in different interstellar environments, Fan et al. (2017) analyzed how the pattern of the equivalent widths (EWs) of DIBs —normalized by the corresponding extinction— changes with the molecular fraction $f_{\rm H2}$.
@@
The first observed DIBs, and those seen to have the strongest signal (Merrill 1930), the DIBs $\lambda 5780$, $\lambda 5797$, and $\lambda 6614$, as well as their interrelations, have been extensively studied and discussed in the literature. Owing to the proximity of the DIBs $\lambda 5780$ and $\lambda 5797$, they are usually investigated as a pair. The well-known “$\sigma-\zeta$” effect was found from the varying strength ratio of these two prominent DIBs (e.g., Krelowski & Westerlund 1988; Vos et al. 2011; Kos & Zwitter 2013). These works indicated that the EWs of the DIB $\lambda 5780$ in the “$\zeta-$type” sight lines tend to be weaker than those in the “$\sigma-$type” sight lines, while the corresponding EWs of the DIB $\lambda 5797$ are similar. Upon discovering that the relationship between DIBs and reddening is linear, but with considerable dispersion (including the DIBs $\lambda 5780$, $\lambda 5797$, and $\lambda 6614$; Raimond et al. 2012; Puspitarini et al. 2013), and that there exists a “skin effect” (Snow & Cohen 1974; Herbig 1995), the correlations with other ISM tracers, such as OH, CH, CO, H  I and $\rm H_{2}$, have been explored (Weselak et al. 2004; Friedman et al. 2011; Lan et al. 2015; Fan et al. 2017; Weselak 2019). It is now well established from those studies that the DIB $\lambda 5797$ is more closely related to the column density of $\rm H_{2}$ than DIB $\lambda 5780$, while this latter shows a stronger correlation with that of H  I . In addition, through studies of the correlations between DIBs, an almost perfectly correlated DIB pair has been identified, namely DIB $\lambda 6196$-$\lambda 6614$ (Cami et al. 1997; Moutou et al. 1999; Galazutdinov et al. 2002; McCall et al. 2010; Bailey et al. 2016), as have some relatively well correlated DIB pairs, such as DIB $\lambda 5780$-$\lambda 6614$ and DIB $\lambda 5797$-$\lambda 5850$ (Friedman et al. 2011; Smith et al. 2021; Fan et al. 2022), and even some anticorrelated DIB pairs, such as DIB $\lambda 4984$-$\lambda 7559$ and $\lambda 5418$-$\lambda 7562$ (Fan et al. 2022), which can help to group the DIBs into specific DIB families. Furthermore, in high-resolution spectra, the longward blue wing in the DIB $\lambda 5780$ profile (Snell & vanden Bout 1981; Danks & Lambert 1976), the asymmetry profile in the DIB $\lambda 5797$ (MacIsaac et al. 2022), and the triple-peak fine structure in the DIB $\lambda 6614$ profile (Herbig & Soderblom 1982; Galazutdinov et al. 2008), all suggest that their carriers are complex macromolecules.
@@
However, the majority of the aforementioned studies are mainly based on the high-resolution spectra of early-type stars without contamination from the stellar lines. It is challenging to carry out larger-scale and more statistically significant analyses as a result of the limited sight lines toward the early-type stars. Kos et al. (2013) introduced a best neighbor matching method (BNM) to extract the DIB $\lambda 8621$ from approximately 500,000 spectra of late-type stars in the Radial Velocity Experiment survey (RAVE; Steinmetz et al. 2006) and provided a statistical ratio between DIB $\lambda 8621$ and extinction. Although the objects observed in the Sloan Digital Sky Survey (SDSS; Yanny et al. 2009) do not concentrate on the Galactic plane where the reddening is high, Lan et al. (2015) still managed to measure over 20 DIBs from the spectra of cool stars by stacking the spectra with similar extinction, and mapped out the distribution of DIBs at high latitudes in the Milky Way. Later on, using the cool stars from the Apache Point Observatory Galactic Evolution Experiment (APOGEE; Majewski et al. 2017), Zasowski et al. (2015) presented a projected Galactic map of the single DIBs at 1.5273 $\mu$m and were the first to estimate its precise central wavelength based on kinematics information. Similarly, Gaia Collaboration et al. (2023b) and Gaia Collaboration et al. (2023a) further elaborated the properties of the DIBs $\lambda 8621$ and $\lambda 8648$ in the $Gaia$ radial velocity spectrometer (RVS) spectra (Cropper et al. 2018; Seabroke et al. 2021) of both cool and hot stars. Furthermore, Cox et al. (2024) constructed a 3D Galactic map tracing the DIB $\lambda 8621$ based on the measurements of Gaia Collaboration et al. (2023b). Using the information provided by the GALactic Archaeology with HERMES (GALAH; Buder et al. 2021) for approximately 872,000 cool stars, brand new DIBs were found in the residuals of these stellar spectra (Vogrinčič et al. 2023). Moreover, Saydjari et al. (2023) and Zhao et al. (2024) recently employed the methods based on machine learning to improve measurements of DIBs from the spectra of cool stars. To measure DIBs from cool stellar spectra, the stellar lines must be removed from spectra in order to extract the ISM composition. A critical distinction exists between the methods used by authors to obtain the stellar lines. On the one hand, studies such as those of Chen et al. (2013), Puspitarini et al. (2015), Zhao et al. (2021a), and Gaia Collaboration et al. (2023b) use theoretical methods to synthesize the stellar lines. On the other hand, observed spectra with low or no extinction are used to pairwise match the stellar components (e.g., Cordiner et al. 2008a, b; Kos & Zwitter 2013; Kos et al. 2014; Zasowski et al. 2015 and Gaia Collaboration et al. 2023a), which is also the approach adopted in the present study.
@@
In addition to the works of Yuan & Liu (2012) and Lan et al. (2015), which are based on low-resolution spectra (R $\approx$ 2000) obtained by SDSS, certain teams have also made efforts to glean hidden information from low-resolution spectra. Using spectra with a resolving power of R = 3300, Cordiner et al. (2008a) investigated the DIBs $\lambda 5780$, $\lambda 5797$, $\lambda 6203$, $\lambda 6283$, and $\lambda 6614$ in the ISM of M31. These authors then used the spectra with similar resolution to extend their work to the DIBs in the ISM of M33 (Cordiner et al. 2008b), and carried out the first survey of DIBs observed in the spectra of B-type supergiants in M31 (Cordiner et al. 2011). By means of the DIBs $\lambda 5780$ and $\lambda 5797$ measured from spectra (Farhang et al. 2015a) with R = $\sim$ 2000 in the northern hemisphere, Farhang et al. (2015b) probed the Local Bubble and its surroundings. Later, combining the DIB measurements derived from the spectra with a resolution of 5500 in the southern observations, they presented a 3D map of the hot Local Bubble (Farhang et al. 2019). Although the quantity of low-resolution spectra used in these studies has risen tremendously compared to high-resolution spectra, it is still insufficient to support a broader perspective for investigating the properties of DIBs, such as their spatial distribution and kinematics in the Milky Way.
@@
In the present work, we take full advantage of the LAMOST low-resolution survey (LAMOST LRS; Zhao et al. 2012; Deng et al. 2012; Luo et al. 2015). The survey almost covers the entire northern sky from the declination of -10 to nearly 90, with particularly dense coverage of the ISM-rich Galactic plane. So far, a few million spectra have been obtained for cool stars across the optical wavelength range, which allow us to detect several DIBs simultaneously. We selected the DIBs $\lambda 5780$, $\lambda 5797$, and $\lambda 6614$ as our subjects, because their higher intensities compared to the other DIBs (Fan et al. 2019; Vogrinčič et al. 2023) mean that we are more likely to be able to measure them in the individual spectra of cool stars. Although the DIBs $\lambda 4430$ and $\lambda 6283$ are also strong, they are not included in this work because of the complexity of the spectral region in which they are found. For the DIB $\lambda 4430$, it is hard to determine the continuum level due to the adjacent strong hydrogen Balmer lines and its own broad profile. For the DIB $\lambda 6283$, the O2 telluric band wraps around it (Herbig 1975; Jenniskens & Desert 1994) and is not cleanly removed in the LAMOST spectra. As a consequence, we focus on the DIBs $\lambda 5780$, $\lambda 5797$, and $\lambda 6614$, build the largest sample of measurements of these three DIBs to date, and conduct a comprehensive and statistically robust analysis of their properties.
@@
The remainder of this paper is organized as follows. In Sect. 2, we describe the data used in this work. The pipeline for measuring the DIBs $\lambda 5780$, $\lambda 5797$, and $\lambda 6614$ in the spectra of cool stars is detailed in Sect. 3. In Sect. 4, we control the quality of our measurements and carry out a series of validation tests to ensure the reliability of the measurements. We present and discuss our results in Sect. 5. Section 6 introduces the code and data availability. Finally, we summarize our findings in Sect. 7.
@@ 2 Data
LAMOST (the Large Sky Area Multi-Object-fiber Spectroscopic Telescope; Zhao et al. 2012; Deng et al. 2012; Luo et al. 2015) is a reflecting Schmidt telescope with a $5^{\circ}$ field-of-view and an effective aperture of 4 meters. Thanks to 4000 fibers averagely distributed among 16 spectrometers (i.e., each one accepts 250 fibers), LAMOST is able to simultaneously observe 4000 objects in theory. There are two resolving modes for LAMOST: one is the low resolution (R = 1800 at 5500 Å) and the other is the medium resolution (R = 7500 at 5163 Å and 6593 Å). When switching to the low-resolution mode employed here, the designed optical band covers the range from 3700 Å to 9000 Å.
The data used in this work are from the data release 10 of LAMOST low-resolution spectroscopic survey (LAMOST LRS DR10)  [^2] https://www.lamost.org/dr10.. A total of 11,817,430 spectra are published in this release, which contains 11,473,644 stellar spectra, 263,444 galaxy spectra, and 80,342 quasar spectra. All the published spectra are reduced by wavelength calibration, sky background subtraction, telluric correction, and relative flux calibration (Zhao et al. 2012; Luo et al. 2015). Note that the telluric absorption is negligible in the regions of the DIBs $\lambda 5780$, $\lambda 5797$, and $\lambda 6614$ investigated in this work (see Fig. 16 in Matheson et al. 2000 and Fig. 2 in Kimeswenger et al. 2015 for more details). All the LAMOST targets are cross-matched with $Gaia$ DR3 sources (Gaia Collaboration et al. 2023c) within a radius of 3 arcsec, and the $Gaia$ IDs, namely gaia_source_id, are included in the official LAMOST DR10 catalog  [^3] https://www.lamost.org/dr10/v1.0/catalogue. The stellar spectra we focus on are corresponding to $\sim$2,000,000 unique stars and the majority of them concentrate on the Galactic plane. For the cool stars, that is, the F, G, K, and late A-type stars, the LAMOST stellar parameter pipeline (LASP) provides the effective temperature ($T_{\rm eff}$), surface gravity ($\log g$), metallicity abundance ([Fe/H]), and radial velocity (RV) with the precision of 100 K, 0.19 dex, 0.14 dex, and 6 $\rm km\ s^{-1}$, respectively  [^4] Here we refer reader to the note of LAMOST DR10 via https://www.lamost.org/dr10/v1.0/doc/release-note.. In addition, the continuum-normalized flux are calculated for these spectra with stellar parameters (Song et al. 2012).
Benefiting from the estimations of geometric distances from parallaxes in $Gaia$ DR3 (Bailer-Jones et al. 2021), the relatively reliable distance of an object in LAMOST can be simply derived. Combined with the three-dimensional dust reddening map (version name “bayestar2019”, Green et al. 2019), once the celestial coordinate and astrometric distance of an object is given, the fine-grained extinction $\rm E(B-V)$  [^5] A recalibration factor of 0.884 (Schlafly & Finkbeiner 2011) is applied for $\rm E(B-V)$. can be calculated at the sight line of each object.